Aggregator

CREW CHIEF SIGN IN

1 day 12 hours ago
Log Number: 4468770Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 02-05-26
Time : 23:03:32
Day : Thursday
Shift : Owl

Crew Chief : Price
Program Deputy: Mock
Team : D
Arm : RADCON
Designated SSO: Howard
LERF Operator:
Operators : Howard

This stamp certifies that the above Crew Chief has read/reviewed/received the following:
kprice

Ops-PR kicked back as "undeliverable"

1 day 13 hours ago
Log Number: 4468769Logbook(s): Entry Makers: eformanBody: 

Filed an ops-pr (elog 4468733) that returned the following message to my email inbox:

----------------------------
This is the mail system at host sm5.jlab.org.

I'm sorry to have to inform you that your message could not
be delivered to one or more recipients. It's attached below.

For further assistance, please send mail to postmaster.

If you do so, please include this problem report.

eforman

Injector steering from FC1 to FC2

1 day 13 hours ago
Log Number: 4468768Logbook(s): Entry Makers: eformanBody: 

Steered injector from FC1 to FC2.

Centering through buncher proved to be a bit tricky because I couldn't see the dithering (that should be a good thing, right?) I steered off of center just so I could see what dithering looks like off-center (I wanted to see what "bad" looked like before I declared it "good"). However, using the recommended correctors, I couldn't seem to make the centering bad enough to see the dithering unless I cranked the step size up to +/- 45 degrees.

eforman

0L06 119.6 MeV Energy Measurement

1 day 13 hours ago
Log Number: 4468767Logbook(s): Entry Makers: bschaumlBody: 

Checked energy to the 4D00 viewer with MBF0L06 set for 119.6 MeV. Phases needed to be tweaked a smidge, but made up most of the low energy. I did have to add 0.1 MV to each lock zone cavity to get it right on the money... not sure if that is proper or not.

bschauml

MOLLER Install Update 02/05/2026

1 day 14 hours ago
Log Number: 4468766Logbook(s): Tags: Entry Makers: jonesgaskelldwinesdougberdnikciprianBody: 

Hall A techs continued moving, of MPS bunker blocks into Hall. Hall a tech completed needed documentation, for upcoming critical lift of pion donut detector pipe. Survey and alignment completed documentation for work of aligning dump pipe/SAM pipe.

lhurt

LEMi

1 day 16 hours ago
Log Number: 4468761Logbook(s): Tags: Body: 

LEMi 3.12 successfully applied the gradient profile to the injector with an elapsed time of 00:00:02.

bschauml

Updated Injector MMS

1 day 17 hours ago
Log Number: 4468754Logbook(s): Entry Makers: bschaumlBody: 

Will be working on high energy setup for upcoming ILD beam studies so set MMS appropriately for running LEMi.

bschauml

COMMENT ON Install MOLLER ADCs, VME modules and connect BCM signal cables in Injector Service Building

1 day 17 hours ago
Log Number: 4468753Logbook(s): Body: 

Installed 4 (out 6 Moller ADCs). Verified the new switch in rack IN03B25 has Hall-A DAQ subnet, and ADCs and VME CPU are powered ON and booted via pxe boot using Hall-A network (similar to the ADCs in Hall-A). The crateHelicity Decoder board is also installed in the VME crate and all helicity related fibers are connected to it.

jpoudel

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Blackshaw
MCC Operators Thomas
Program Deputy Mock
Announcements

Access Status

Recent PSS History

recent pss history graph